# Prompt Design and Engineering: Introduction and Advanced Methods

### 24 Jan 2024 v1

@@ 3 Augmenting LLMs through external knowledge - RAG
One of the main limitations of pre-trained LLMs is their lack of up-to-date knowledge or access to private or use-case-specific information. This is where retrieval augmented generation (RAG) comes into the picture [13]. RAG, illustrated in figure 12, involves extracting a query from the input prompt and using that query to retrieve relevant information from an external knowledge source (e.g. a search engine or a knowledge graph, see figure 13 ). The relevant information is then added to the original prompt and fed to the LLM in order for the model to generate the final response.
@@ 3.1 RAG-aware prompting techniques
Because of the importance of RAG to build advanced LLM systems, several RAG-aware prompting techniques have been developed recently. One such technique is Forward-looking Active Retrieval Augmented Generation (FLARE)
Forward-looking Active Retrieval Augmented Generation (FLARE)[16] enhances the capabilities of Large Language Models (LLMs) by iteratively combining prediction and information retrieval. FLARE represents an evolution in the use of retrieval-augmented generation, aimed at improving the accuracy and relevance of LLM responses.
FLARE involves an iterative process where the LLM actively predicts upcoming content and uses these predictions as queries to retrieve relevant information. This method contrasts with traditional retrieval-augmented models that typically retrieve information once and then proceed with generation. In FLARE, this process is dynamic and ongoing throughout the generation phase. In FLARE, each sentence or segment generated by the LLM is evaluated for confidence. If the confidence level is below a certain threshold, the model uses the generated content as a query to retrieve relevant information, which is then used to regenerate or refine the sentence. This iterative process ensures that each part of the response is informed by the most relevant and current information available.
For more details on RAG and its relevant works, we refer the readers to this survey of retrieval augmented generations [17].
@@ 4 LLM Agents
The idea of AI agents has been well-explored in the history of AI. An agent is typically an autonomous entity that can perceive the environment using its sensors, make a judgment based on the state it currently is, and accordingly act based on the actions that are available to it.
In the context of LLMs, an agent refers to a system based on a specialized instantiation of an (augmented) LLM that is capable of performing specific tasks autonomously. These agents are designed to interact with users and environment to make decisions based on the input and the intended goal of the interaction. Agents are based on LLMs equipped with the ability to access and use tools, and to make decisions based on the given input. They are designed to handle tasks that require a degree of autonomy and decision-making, typically beyond simple response generation.
The functionalities of an LLM-based agent involves include:
Tool Access and Utilization: Agents have the capability to access external tools and services, and to utilize these resources effectively to accomplish tasks.
Decision Making: They can make decisions based on the input, context, and the tools available to them, often employing complex reasoning processes.
As an example, an LLM that has access to a function (or an API) such as weather API, can answer any question related to the weather of the specific place. In other words, it can use APIs to solve problems. Furthermore, if that LLM has access to an API that allows to make purchases, a purchasing agent can be built to not only have capabilities to read information from the external world, but also act on it [18].
@@ 4.1 Prompt engineering techniques for agents
Like RAG and Tools, prompt engineering techniques that specifically address the needs of LLM-based agents have been developed. Three such examples are Reasoning without Observation (ReWOO), Reason and Act (ReAct), and Dialog-Enabled Resolving Agents (DERA).
@@ 4.1.1 Reasoning without Observation (ReWOO)
Reasoning without Observation (ReWOO)[19] aims to decouple reasoning from direct observations. ReWOO operates by enabling LLMs to formulate comprehensive reasoning plans or meta-plans without immediate reliance on external data or tools. This approach allows the model to create a structured framework for reasoning that can be executed once the necessary data or observations are available.
In ReWOO, the LLM initially develops a plan or a series of steps that outline how to approach and solve a given problem. This meta-planning phase is crucial as it sets the stage for how the model will process information once it becomes available. The execution phase then involves integrating actual data or observations into the pre-formulated plan, leading to a more coherent and contextually relevant response. ReWOO offers significant advantages in terms of token efficiency and robustness to tool failure. It enables LLMs to handle tasks where immediate access to external data is not available, relying instead on a well-structured reasoning framework. This method is particularly advantageous in scenarios where data retrieval is costly, slow, or uncertain, allowing the LLM to maintain a high level of performance and reliability.
@@ 4.1.2 Reason and Act (ReAct)
Reason and Act (ReAct)[20] prompts LLMs to generate not only verbal reasoning but also actionable steps, enhancing the model’s dynamic problem-solving capabilities. ReAct is grounded in the principle of integrating reasoning with action. In this approach, the LLM is prompted to alternate between generating reasoning traces (explanations) and actions (steps or commands) in an interleaved manner. This approach allows the model to dynamically reason about a problem and propose concrete actions simultaneously.
@@ 4.2 Dialog-Enabled Resolving Agents (DERA)
Dialog-Enabled Resolving Agents (DERA)[21] involves the creation of specialized AI agents that can engage in dialogue, resolve queries, and make decisions based on interactive exchanges. DERA is structured around the idea of utilizing multiple agents within a dialog context, each with specific roles and functions. These agents can include Researchers, who gather and analyze information, and Deciders, who make final judgments based on the information provided. This division of roles allows for a more organized and efficient approach to problem-solving and decision-making.
DERA is particularly advantageous in scenarios requiring complex decision-making and problem-solving, such as in medical diagnostics or customer service. The collaborative and interactive nature of DERA agents allows them to handle intricate queries with a level of depth and nuance that single-agent systems might struggle with. Moreover, this approach aligns closely with human decision-making processes, making the AI’s reasoning and conclusions more relatable and trustworthy.
@@ 5 Prompt Engineering Tools and Frameworks
The development of advanced prompt engineering techniques has been accompanied by the emergence of various tools and frameworks designed to facilitate the implementation and optimization of these methods. Below is an overview of some of the prominent tools and frameworks in this domain.
Langchain222 is a widely recognized prompt engineering toolkit. Initially focused on supporting Chains, it has evolved to support Agents and various Tools for diverse functionalities, ranging from memory handling to web browsing.
Developed by Microsoft, Semantic Kernel333 is a toolkit designed for skills and planning in C# and Python. It now encompasses capabilities beyond its initial scope, including chaining, indexing, memory access, and plugin development.
Guidance444, another library from Microsoft, utilizes a templating language to support a variety of prompt engineering techniques. Being a more recent addition to the field, it offers modern solutions and supports the methods discussed in this guide.
Developed by NVidia, Nemo Guardrails555 is a tool aimed at building Rails to ensure LLMs operate within desired parameters and guidelines.
LlamaIndex666 focuses on managing the data that feeds into LLM applications, primarily through data connectors and tools.
From Intel, FastRAG777 includes not only the basic Retrieval Augmented Generation (RAG) approach but also more advanced implementations of RAG, aligning with techniques discussed in this guide.
Auto-GPT888 is a popular tool for designing LLM agents, facilitating the creation and management of sophisticated AI agents. More recently, textbfAutoGen by Microsoft has gained popularity for designing agents and multi-agent systems.
In summary, these tools and frameworks play a crucial role in the practical application and advancement of prompt engineering, offering solutions that range from basic prompt management to complex AI agent design and operation.
@@ 6 Conclusion
In just a few months prompt design and engineering has become an extremely important aspect of designing LLM and Generative AI systems. In this paper, we have introduced the basics as well as more advanced methods. The space is rapidly evolving and we should expect to see a lot of innovation to happen very quickly. This introduction and guide should still serve as a good guide with an early historical perspective.

### 08 Feb 2024 v3

@@ 5 Augmenting LLMs through External Knowledge - RAG
In addressing the constraints of pre-trained Large Language Models (LLMs), particularly their limitations in accessing real-time or domain-specific information, Retrieval Augmented Generation (RAG) emerges as a pivotal innovation. RAG extends LLMs by dynamically incorporating external knowledge, thereby enriching the model’s responses with up-to-date or specialized information not contained within its initial training data.
RAG operates by formulating queries from input prompts and leveraging these queries to fetch pertinent information from diverse sources, such as search engines (see figure 22) or knowledge graphs(see figure 23). This retrieved content is seamlessly integrated into the LLM’s workflow, significantly augmenting its ability to generate informed and contextually relevant responses.
@@ 5.1 RAG-aware Prompting Techniques
The advent of RAG has spurred the development of sophisticated prompting techniques designed to leverage its capabilities fully. Among these, Forward-looking Active Retrieval Augmented Generation (FLARE) stands out for its innovative approach to enhancing LLM performance.
FLARE iteratively enhances LLM outputs by predicting potential content and using these predictions to guide information retrieval. Unlike traditional RAG models, which typically perform a single retrieval step before generation, FLARE engages in a continuous, dynamic retrieval process, ensuring that each segment of the generated content is supported by the most relevant external information.
This process is characterized by an evaluation of confidence levels for each generated segment. When the confidence falls below a predefined threshold, FLARE prompts the LLM to use the content as a query for additional information retrieval, thereby refining the response with updated or more relevant data.
For a comprehensive understanding of RAG, FLARE, and related methodologies, readers are encouraged to consult the survey on retrieval augmented generation models, which provides an in-depth analysis of their evolution, applications, and impact on the field of LLMs [18].
@@ 6 LLM Agents
The concept of AI agents, autonomous entities that perceive, decide, and act within their environments, has evolved significantly with the advent of Large Language Models (LLMs). LLM-based agents represent a specialized instantiation of augmented LLMs, designed to perform complex tasks autonomously, often surpassing simple response generation by incorporating decision-making and tool utilization capabilities.
LLM agents can access external tools and services, leveraging them to complete tasks, and making informed decisions based on contextual input and predefined goals. Such agents can, for instance, interact with APIs to fetch weather information or execute purchases, thereby acting on the external world as well as interpreting it.
@@ 6.1 Prompt Engineering Techniques for Agents
The integration of LLMs into agent frameworks has led to the development of novel prompt engineering techniques, including Reasoning without Observation (ReWOO), Reason and Act (ReAct), and Dialog-Enabled Resolving Agents (DERA), each tailored to enhance the autonomous functionality of LLM-based agents.
@@ 6.1.1 Reasoning without Observation (ReWOO)
ReWOO enables LLMs to construct reasoning plans without immediate access to external data, relying instead on a structured reasoning framework that can be executed once relevant data becomes available (see figure 25). This approach is particularly useful in scenarios where data retrieval is costly or uncertain, allowing LLMs to maintain efficiency and reliability.
@@ 6.1.2 Reason and Act (ReAct)
ReAct (see figure 26) enhances LLMs’ problem-solving capabilities by interleaving reasoning traces with actionable steps, facilitating a dynamic approach to task resolution where reasoning and action are closely integrated.
@@ 6.2 Dialog-Enabled Resolving Agents (DERA)
DERA (see figure 27) introduces a collaborative agent framework where multiple agents, each with specific roles, engage in dialogue to resolve queries and make decisions. This multi-agent approach enables handling complex queries with depth and nuance, closely mirroring human decision-making processes.
The development of LLM-based agents and associated prompt engineering techniques represents a significant leap forward in AI, promising to enhance the autonomy, decision-making, and interactive capabilities of LLMs across a wide range of applications.
@@ 7 Prompt Engineering Tools and Frameworks
The proliferation of advanced prompt engineering techniques has catalyzed the development of an array of tools and frameworks, each designed to streamline the implementation and enhance the capabilities of these methodologies. These resources are pivotal in bridging the gap between theoretical approaches and practical applications, enabling researchers and practitioners to leverage prompt engineering more effectively.
Langchain has emerged as a cornerstone in the prompt engineering toolkit landscape, initially focusing on Chains but expanding to support a broader range of functionalities including Agents and web browsing capabilities. Its comprehensive suite of features makes it an invaluable resource for developing complex LLM applications.
Semantic Kernel, by Microsoft, offers a robust toolkit for skill development and planning, extending its utility to include chaining, indexing, and memory access. Its versatility in supporting multiple programming languages enhances its appeal to a wide user base.
The Guidance library, also from Microsoft, introduces a modern templating language tailored for prompt engineering, offering solutions that are aligned with the latest advancements in the field. Its focus on modern techniques makes it a go-to resource for cutting-edge prompt engineering applications.
Nemo Guardrails by NVidia is specifically designed to construct Rails, ensuring that LLMs operate within predefined guidelines, thereby enhancing the safety and reliability of LLM outputs.
LlamaIndex specializes in data management for LLM applications, providing essential tools for handling the influx of data that these models require, streamlining the data integration process.
From Intel, FastRAG extends the basic RAG approach with advanced implementations, aligning closely with the sophisticated techniques discussed in this guide, and offering optimized solutions for retrieval-augmented tasks.
Auto-GPT stands out for its focus on designing LLM agents, simplifying the development of complex AI agents with its user-friendly interface and comprehensive features. Similarly, AutoGen by Microsoft has gained traction for its capabilities in agent and multi-agent system design, further enriching the ecosystem of tools available for prompt engineering.
These tools and frameworks are instrumental in the ongoing evolution of prompt engineering, offering a range of solutions from foundational prompt management to the construction of intricate AI agents. As the field continues to expand, the development of new tools and the enhancement of existing ones will remain critical in unlocking the full potential of LLMs in a variety of applications.
@@ 8 Conclusion
Prompt design and engineering will only become more critical as LLMs and generative AI evolve. We discussed foundations and cutting-edge approaches such as Retrieval Augmented Generation (RAG) – essential tools for the next wave of intelligent applications. As prompt design and engineering rapidly progress, resources like this will offer a historical lens on early techniques. Remember, innovations like Automatic Prompt Engineering (APE) covered here could become standard practice in the years to come. Be part of shaping the trajectory of these exciting developments!